how could you make a 7.8% aqueous solution of glucose using 5.0g of glucose?

1
Expert's answer
2022-03-22T14:47:03-0400

Let x be mass of water


"\omega(glucose)=\dfrac{mass\ of\ glucose}{mass\ of\ solution}=\dfrac{5.0\ g}{5.0+x\ g}=0.078"


"5.0*0.078+0.078x=5.0"


"0.078x=5.0-5.0*0.078=4.61"


"x=\dfrac{4.61}{0.078}=59.1\ g"


Answer: add 59.1 g of water